if y=(x+(x^2+m^2)^1/2)^n then dy/dx
1
Expert's answer
2012-10-12T10:48:12-0400
y'=(x+(x^2+m^2)^1/2)^n=(n(x+(x^2+m^2)^1/2))^(n-1))(x+(x^2+m^2)^1/2)'=
(n(x+(x^2+m^2)^1/2))^(n-1))(1+((x^2+m^2)^1/2)'=(n(x+(x^2+m^2)^1/2))^(n-1))(1+2x/2((x^2+m^2)^1/2)=(n(x+(x^2+m^2)^1/2))^(n-1))(1+x/((x^2+m^2)^1/2)